PARTI DETERMINATION OF x in KCIO DATA Unknown Identification Letter or Number Mass of test tube + MnO2 1.0979 Mass of test
Answer #1

The ratio is calculated:

Ratio = n O2 / n KCl = 0.01472 / 0.007238 = 2

The ratio has the balanced reaction:

KClO4 = KCl + 2 O2

b = 1

c = 2

a = 1

x = 4

The percentage of experimental oxygen is calculated:

% O exp = 0.4709 * 100 / 1.0104 = 46.61%

% O theoretical = 4 * 16 * 100 / 138.55 = 46.19%

The error percentage is calculated:

% error = (46.61 - 46.19) * 100 / 46.19 = 1%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.